Use the following atomic weights & quantities to calculate overall % yield of diphenylacetylene. Remember that your yield must be based on the limiting reagent, and that grams must be converted to moles. For problems not involving a one to one stoichiometry ot may be easier to calculate the theoretical yield in grams from both reagents to determine which is truly the limiting reagent. Filling out most of the chart will help. Give ONLY two significant digits in your answer. If after rounding the answer is a whole # , do not include a decimal point.

OctoberDiphenylacetylene Report Question 9 Not yet answered yield must be based on the limiting reagent, and that grams must be converted to moles. For problems not involving a one Marked out of Use the following atomic weights and quantities to calculate the overall % yield of diphenylacetylene Remember that your to one stoichiometry it may be easier to calculate the theoretical yield in grams from both reagents to determine which is truly the limiting reagent. Filling out most of the chart will help. Give only two significant digits in your answer. If ater rounding the answer is a whole number, do not include a decimal point. C 12, H 1, Br 80, K 39, O 16 P Flag question stilbene dibromide +potassium hydroxide diphenylacetylene formula formula weight grams moles 68.063 5.720 8.908 theoretical yield (g) % yield Answer
Answer #1

C14H12Br2+2KOH-------C14H10+2H2O+2KBr

1 mol. 2mol. 1 mol. 2mol. 2 mol

68.063g. 5.720g. 8.908g

.200mol. .102mol.   

.200/1. .102/2

.051 mol

So KOH will be the limited agent

Now, 2 mol KOH gives= 1 mol C14H10

1 molKOH= 1/2 mol

.051 mol KOH = 1/2*.051

= .0255mol

=.0255*178 g

=4.539g

% yield = 4.539/8.908*100

=50.9542%